OpenStudy (anonymous):

what is the solution to the equation sqrt5^8n= 125^(n+3)?

OpenStudy (nikvist):

\[(\sqrt{5})^{8n}=125^{n+3}\]\[(5^{1/2})^{8n}=(5^3)^{n+3}\]\[4n=3n+9\]\[n=9\]